OpenStudy (anonymous):

Please help! Evaluate each of the following without using a calculator or computer. Show your work. a. sin(arcsin(1/2) + arcos(1/2)) b. sec(arccos(-√3/2)) c. arctan(tan(3π/4)) d. tan(arctan(3π/4))

OpenStudy (anonymous):

\[\sin\left[\sin^{-1}{1\over2}+\cos^{-1}{1\over2}\right]\\ \quad=\sin\left[{\pi\over6}+{\pi\over3}\right]=\sin\left(\pi\over2\right)=? \]

OpenStudy (anonymous):

give others a shot in the same way